# Synth: Boosting Visual-Language Models with Synthetic Captions and Image Embeddings

### v1

@@ 4.1.2 Training details
For both the image generator and the VLM we use a pre-trained and frozen VQ-GAN (see A.1.1 for details on the network). The images are input at a resolution of 256x256, with a patch size of 16, which results in a sequence length of 256 tokens (16x16). Both models use a pre-trained and frozen Chinchilla 400m (Hoffmann et al., 2022) as LLM. The VLM has a perceiver resampler (Alayrac et al., 2022; Jaegle et al., 2021) with 6 layers and 64 latent input queries that cross-attend to the image embeddings. There are cross attention layers between each layer in the LLM and the output of perceiver resampler similar to the original Flamingo architecture (please refer to A.1.4 for extra details).
We use ViT-B (Dosovitskiy et al., 2020) for the main text-to-image model in the image generator (a transformer with 24 layers, embedding dimension of 769, 12 heads and an intermediate latent size of 3072, dropout rate of 0.1). We use a maximum text length of 64 tokens with guidance scale of 4.0 with 24 refining steps and sample choice temperature of 32.5. All the models are trained with AdamW (Loshchilov & Hutter, 2017) with a learning rate of 0.0001 with 5000 warmup steps, a batch size of 512. For finetuning settings (on COCO), we use a learning rate of 0.00001. Section A.1.3 reports extra details. We pre-train the image generator for 500k steps at which point it has a Fréchet inception distance (FID) (Heusel et al., 2017) of 17.1 on MS-COCO test pairs. Our VLMs training experiments all run for 500k steps. Our models are all implemented in Jax and trained on 256 TPUs.
@@ 4.2.1 Synthetic Images
To assess the effectiveness of synthetic images in VLM training, we conduct a study where human-written captions from existing datasets are paired with synthetic images generated by a text-to-image model. We train a VLM on these synthetic image-caption pairs and evaluate its performance against a model trained on the original real image-caption pairs. This comparison allows us to investigate whether synthetic images can effectively substitute for real images in the context of VLM training.
Specifically, we generate synthetic images for the LTIP and WebLI datasets using their annotations. This provides us with (Real Text, Synthetic Image) pairs to train a VLM with (Synth in Table 2). We compare this model to the models trained on the original (Real Text, Real Image) pairs (Gold in Table 2) and one trained from human-annotated data of Conceptual Caption v2 (CCv2) only, using no data from LTIP and WebLI (Baseline in Table 2). For consistency, all Synth and Gold models are co-trained with CCv2, the same dataset used by our text-to-image generator and the baseline.
| Real | Synth | #Real Data | #Synth Data | Concentration () | Entropy () | CIDEr-COCO () |
| CCv2 | - | 10.1M | - | - | - | 22.1 |
| CCv2 | GenPair | 10.1M | 1M | 57.7% | 3.81 | 25.9 |
| CCv2+WebLI | - | 10.1M+1M | - | 69.8% | 3.43 | 24.4 |
| CCv2+LTIP | - | 10.1M+1M | - | 83.0% | 2.92 | 23.4 |
As shown in Table 2, synthetic images significantly improve baseline model performance by over 25%. Importantly, they are effective for VLM training despite the smaller volume of human-annotated images used. Note that “Gold” represents the upper performance bound when using the full original dataset. For LTIP, synthetic images slightly outperform original images, likely due to increased data diversity. Because original pairs remain static during training, synthetic images introduce beneficial image augmentation (Azizi et al., 2023).
These results suggest the potential of generative models to enhance VLM training, even with limited real image data. While previous work (Azizi et al., 2023; Fan et al., 2023) demonstrated the use of synthetic images for single-object image classification (i.e., ImageNet), our findings show that VLMs can be successfully trained with synthetic datasets containing complex, multi-object scenes and detailed captions. This highlights the potential of synthetic images to move beyond simple classification and significantly improve VLM performance in diverse, challenging scenarios.
@@ 4.2.2 Synthetic text and image pairs
To further demonstrate the efficacy of using synthetic data for VLM training, we conducted additional experiments where the entire dataset including captions and their corresponding images are generated synthetically by an LLM and the image generator (see Section 3.3 for the explanation of how these caption where generate and Figure 2 for some examples).
Comparing the first two rows in Table 3 (baseline vs. GenPair) shows that adding even a small fraction (1M) of purely synthetic image-caption data (GenPair) significantly improves performance (17.2% increase in CIDEr score on MS-COCO compared to the baseline model trained only on real data). Interestingly, the same improvement is not achieved by sampling an additional 1M data points from real datasets like WebLI or LTIP (rows 3 and 4). To investigate this, we assessed the semantic diversity and balance of each caption set. First, we used our language model to generate embeddings for the captions. Then, we employed k-means clustering to analyze these embeddings (see Appendix A.3 for details). Figure 5 visualizes the cluster distribution, with the x-axis representing the cluster index and the y-axis representing data volume from each dataset within a cluster. GenPair shows a more even distribution across clusters, indicating greater conceptual diversity. It also exhibits less variance in cluster size, suggesting better balance.
Finally, to quantify the semantic concentration within each dataset, Table 3 presents the ”Concentration” metric: the percentage of data belonging to the top-5 most populated clusters. GenPair has the lowest concentration, with only 57.5% of its captions in the top-5 clusters. This indicates lower semantic imbalance in the synthetically generated GenPair and contrasts with the higher imbalance found in the other datasets (69.9% and 83%). Also, the entropy over the full cluster distribution, confirms that GenPair has a more uniform distribution of data across clusters. We postulate that the inherent diversity within GenPair likely contributes to its robust performance on MS-COCO, as models trained on diverse data tend to generalize better. (See Appendix A.3 for extra details on how the cluster analysis was performed and for extra results.)
This evaluation showcases Synth’s ability to utilize diverse text-based data, including purely synthetic ones. Overall, our experiments highlight the significant role that synthetic data can play in enhancing VLM performance by leveraging synthetic image generation and LLM-generated captions. This approach not only overcomes data limitations but also demonstrates the effectiveness of text-only captions in boosting model capabilities, further emphasizing the value of our technique.
@@ 4.2.3 Comparison with related work
Furthermore, we investigated the potential of our approach compared to the state-of-the-art model while taking into account the amount of synthetic or real training data and the number of model parameters. We combined the two datasets with the highest diversity, namely, GenPair and WebLI. Table 4 compares our model’s MS-COCO benchmark performance against related methods: ITIT (Li et al., 2023b) and DC (Li et al., 2023c). Note that, for consistency with these methods, we used Synth finetuned on the MS-COCO training set (see A.2). ITIT employs cycle consistency for VLM improvement, while DC uses Stable Diffusion to generate images based on human-defined heuristics.
@@
Synth and ITIT share similar parameter counts and paired data usage, making them directly comparable. As shown in Table 4, Synth drastically outperforms ITIT on the CIDEr score. While DC achieves the highest raw performance, it requires significantly more parameters, relies on vast amount of real data, and the filtering procedure is heavily based on human heuristics.
| Model | All Params | Real Data | Synth Data | CIDEr (FT) |
| ITIT | 11.2B | 3M | 110M | 103.5 |
| DC | 1.7B | 5.5B | 400M | 133.1 |
| Synth | 632M | 10.1M | 360M | 126.1 |
@@
Synth strikes a balance between performance and efficiency. By leveraging text-only data, it achieves competitive results using 40 times less paired data than DC. This comparison highlights the trade-offs between performance, efficiency, and data requirements. Synth emerges as a compelling solution when resources are limited, achieving impressive results with reduced data and computational overhead.
4.2.4 Data efficiency and augmentation with fully synthetic data
@@
To further analyze the boost in performance afforded by augmentation with fully synthetic data, we characterized the performance of the baseline model (i.e. trained solely on CCv2), and Synth which was additionally trained together with GenPair. Figure 6 visualizes the performance trends for both models. Notably, the light blue curve representing the baseline exhibits a gradual increase in performance with increasing training steps, eventually plateauing at 22.1 CIDER score. In contrast, the purple curve reveals Synth’s steeper performance improvement, achieving comparable performance to the paired only training regime, with roughly 1/3 of the training steps. This highlights the significant data efficiency gain achieved through fully synthetic data augmentation in Synth. Shaded regions surrounding each curve indicate standard deviation across three random seeds, demonstrating the robustness of the observed performance trends. These findings demonstrate Synth’s ability to effectively leverage fully synthetic data, resulting in remarkable performance gains.
@@
Our study also investigated the computational efficiency of Synth. We compared the case where captions are used to generate image embedding versus when captions are rendered into actual images. As shown in Table 5, Synth trained using image embedding consistently demonstrates faster performance, running at 2.08 training steps per second, whereas Synth trained from pixel runs at a slower pace of 1.66 steps/second. This performance advantage stems from Synth’s efficient utilization of text-only data and embedding-based image generation using parallel decoding, which reduces the computational overhead associated with processing pixel-based images as done in the Synth model trained from pixels without affecting the performance.
@@ 5 Limitations
The three main limitations of the current work stem from the restricted quantity of fully synthetic data used, potential biases in the generative model and the limited exploration of text data sources. Firstly, whilst we were able to show a substantial gain when training was augmented with fully synthetic data, our experiments were limited to a relatively low quantity of data (i.e. 1M image-caption pairs). In the future, it will be important to examine whether using orders of magnitude more fully synthetic data (e.g. 700M) will result in performance gains that surpass that of using solely paired data. Secondly, the generative model responsible for text-to-image generation can introduce biases into the synthetic data, reflecting limitations inherent in its training data, architecture, or procedure. A possible future step would involve investigating this issue using various text-to-image models to assess and mitigate potential biases. Thirdly, the current work primarily explores a restricted set of text-only data sources. Further research is necessary to investigate the potential of diverse text sources, including investigating context-specific generation (e.g., medical data), which could be an exciting avenue for future exploration.
By acknowledging and addressing these limitations, future research can leverage the strengths of the Synth approach and enhance its generalizability, robustness, and efficiency for a broader range of visual-language modeling tasks.
| Model | Step/Sec | CIDEr(zs) |
| Synth- Image Embedding | 2.08 | 25.9 |
| Synth- Pixels | 1.66 | 26.0 |
@@ 6 Conclusions
Our study presents a novel approach for generating synthetic image-text pairs to enhance the training of visual language models. By harnessing the capabilities of large language models and text-to-image generation, we effectively address the limitations of manual image labeling. Our results conclusively demonstrate several significant findings:
Improved VLM Performance: The visual language model trained on our synthetic and human annotated datasets exhibits a marked improvement in image captioning tasks compared to a baseline trained exclusively on human-annotated data. This underscores the potential of our method to augment VLM capabilities efficiently. This is highly advantageous in scenarios where data acquisition and annotation are resource-intensive.
Data Efficiency: Our approach yields comparable performance while utilizing only a fraction of human-labeled data.
Customization and Scalability: Our method provides flexibility in generating customized image datasets tailored to specific domains. Additionally, the synthetic data generation process is scalable, making it suitable for supporting large-scale VLM development.
This research introduces a novel technique that has the potential to transform visual language model training. Our findings highlight the value of synthetic data generation, paving the way for advancements in numerous fields where visual language understanding is crucial. Further explorations could investigate refining the synthetic data generation process with different objectives in mind.
@@ Acknowledgements
We would like to thank Sander Dieleman, Ali Razavi, and Benigno Uria for their insights on VQGAN. We would also like to thank Aida Nematzadeh, Pinelopi Papalampidi, Han Zhang, Mateusz Malinowski, Valentin De Bortoli, Sahra Ghalebikesabi, Emanuele Bugliarello, Chris Knutsen, and Murray Shanahan for their in depth comments throughout the project.
@@ A.1.1 VQ-GAN details
| Optimizer | Adam (Kingma & Ba, 2014) |
@@
| Decay (Loshchilov & Hutter, 2017) Warmup steps | 10000 |
@@
A.1.2 Optimization details for Synth
@@
| Optimizer | AdamW (Loshchilov & Hutter, 2017) |
@@ A.3 Clustering details
| Real | Synth | #Real Data | #Synth Data | CIDEr-COCO () | Concentration Top-3 () | Concentration Top-5 () | Entropy () |
| CCv2+WebLI | - | 10.1M+1M | - | 24.4 | 58.2% | 69.8% | 3.43 |
| CCv2+LTIP | - | 10.1M+1M | - | 23.4 | 76.5% | 83.0% | 2.92 |
| CCv2 | GenPair | 10.1M | 1M | 25.9 | 40.5% | 57.7% | 3.81 |

### v2

@@ 4.1.2 Training details
For both the image generator and the VLM we use a pre-trained and frozen VQ-GAN (see A.1.1 for details on the network). The images are input at a resolution of 256x256, with a patch size of 16, which results in a sequence length of 256 tokens (16x16). Both models use a pre-trained and frozen Chinchilla 400m (Hoffmann et al., 2022) as LLM. The VLM has a perceiver resampler (Alayrac et al., 2022; Jaegle et al., 2021) with 6 layers and 64 latent input queries that cross-attend to the image embeddings. There are cross attention layers between each layer in the LLM and the output of perceiver resampler similar to (Alayrac et al., 2022) ( refer to A.1.4 for details).
We use ViT-B (Dosovitskiy et al., 2020) for the main text-to-image model in the image generator (see Appendix A.1.3 for full details), with a maximum text length of 64 tokens with guidance scale of 4.0, 24 refining steps and a sample choice temperature of 32.5. All the models are trained with AdamW (Loshchilov and Hutter, 2017) with a learning rate of 0.0001, 5000 warmup steps, and the batch size of 512. For finetuning settings (on COCO), we use a learning rate of 0.00001. Section A.1.3 reports additional details. We pre-train the image generator for 500k steps at which point it has a Fréchet inception distance (FID) (Heusel et al., 2017) of 17.1 on MS-COCO test pairs. Our VLMs training experiments all run for 500k steps. Our models are all implemented in JAX and trained on 256 TPUs.
@@ 4.2 Results
Method | Real | Synth | #Real Data | #Synth Data | CIDEr-COCO () | CIDEr-Flickr-30 () | VQAV2 Accuracy () | OKVQA Accuracy () |
Baseline | CCv2 | - | 10.1M | - | 22.1 | 12.7 | 29.1 | 32.4 |
Synth2 | CCv2 | LTIP | 10.1M | 330M | 28.7 | 16.7 | 33.5 | 34.5 |
CCv2 | WebLI | 10.1M | 350M | 27.6 | 18.1 | 34.1 | 34.7 | |
CCv2 | WebLI+LTIP | 10.1M | 670M | 33.4 | 20.8 | 35.3 | 36.1 | |
Gold | CCv2+LTIP | - | 340.1M | - | 27.6 | 20.9 | 36.7 | 38.1 |
CCv2+WebLI | - | 360.1M | - | 30.7 | 21.1 | 36.2 | 38.5 | |
CCv2+LTIP+WebLI | - | 690.1M | - | 35.3 | 23.4 | 37.9 | 39.4 |
Real | Synth | #Real Data | #Synth Data | Concentration () | Entropy () | CIDEr-COCO () | Flickr-30() | VQAV2 Acc. () | OKVQA Acc. () |
CCv2 | - | 10.1M | - | - | - | 22.1 | 12.7 | 29.1 | 32.4 |
CCv2 | GenPair | 10.1M | 1M | 57.7% | 3.81 | 25.9 | 17.3 | 31.1 | 34.0 |
CCv2+WebLI | - | 10.1M+1M | - | 69.8% | 3.43 | 24.4 | 14.9 | 30.6 | 33.9 |
CCv2+LTIP | - | 10.1M+1M | - | 83.0% | 2.92 | 23.4 | 13.8 | 30.3 | 32.9 |
@@ 4.2.1 Synthetic Images
To first assess the effectiveness of synthetic images compared to the original ones in VLM training, we conduct a study where human-written captions from existing datasets are paired with synthetic images generated by a text-to-image model. We train a VLM on these synthetic image-caption pairs and evaluate its performance against a model trained on the original real image-caption pairs. This comparison allows us to investigate whether synthetic images can effectively substitute for real images in the context of VLM training.
Specifically, we generate synthetic images for the LTIP and WebLI datasets using their annotations. This provides us with (Real Text, Synthetic Image) pairs to train a VLM with (Synth2 in Table 2). We compare this model to the model trained on the human-annotated data of Conceptual Caption v2 (CCv2) only, using no data from LTIP and WebLI (Baseline) and the models trained on the original (Real Text, Real Image) pairs (Gold in Table 2). For consistency, and to reduce model collapse, all Synth2 and Gold models are co-trained with CCv2, the same dataset used by our text-to-image generator and the baseline.
As shown in Table 2, synthetic images significantly improve baseline model performance across the full set of downstream tasks. Importantly, they are effective for VLM training despite the smaller volume of human-annotated images used. Note that “Gold” represents the upper performance bound when using the full original dataset. For LTIP, synthetic images slightly outperform original images, likely due to increased data diversity; while original pairs remain static during training, synthetic images introduce beneficial image augmentation (Azizi et al., 2023).
These results suggest the potential of generative models to enhance VLM training containing complex, multi-object scenes and detailed captions, even with limited real image data.
@@ 4.2.2 Synthetic text and image pairs
To further demonstrate the efficacy of using synthetic data for VLM training, we conducted additional experiments where the entire dataset including captions and their corresponding images are generated synthetically by an LLM and the image generator (see Section 2 for the explanation of how these caption where generate and Figure 1D for some examples).
Comparing the first two rows in Table 3 shows that adding even a small fraction (1M) of purely synthetic image-caption data (GenPair) significantly improves performance across all evaluated domains. Interestingly, sampling an additional 1M data points from real datasets like WebLI or LTIP (rows 3 and 4) yields lower improvement.
To investigate this, we assessed the semantic diversity and balance of each caption set. First, we used our language model to generate embeddings for the synthetic GenPair and the real WebLI and LTIP captions. Then, we employed k-means clustering to analyze these embeddings (see Appendix A.3 for details). Figure 4 shows the cluster distribution among all three datasets, with the x-axis representing the cluster index and the y-axis representing data volume from each dataset within a cluster. Notice how LTIP and WebLI have a large concentration of data only within clusters 19, 18, 1, 16 and 0, while most other clusters are under-populated. In contrast, GenPair is distributed more evenly across different clusters (e.g. clusters 1, 4, 6, 9, 12, 14, 15, 16, 17, 18 are all well populated), suggesting superior balance compared to the other two.
To quantify the mentioned semantic concentration within each dataset we use a “Concentration Metric” as the percentage of data belonging to the top-5 most populated clusters. As reported in Table 3, GenPair has the lowest concentration, with only 57.5% of its captions in the top-5 clusters. This indicates lower semantic imbalance in the synthetically generated GenPair and contrasts with the higher imbalance found in the other datasets (69.9% and 83%). Also, the entropy over the full cluster distribution, confirms that GenPair has a more uniform distribution of data across clusters. We postulate that the inherent diversity within GenPair likely contributes to its robust performance on downstream tasks, as models trained on diverse data tend to generalize better.
This evaluation demonstrates Synth2’s capacity to leverage diverse text data, including synthetic data, for enhanced performance. Our proposed analysis and metrics can inform synthetic data creation, further optimizing its impact. Notably, our findings highlight the significant role of synthetic data, generated via image synthesis and LLM captions, in overcoming data limitations and boosting VLM capabilities. This underscores the effectiveness of text-only captions for model enhancement, emphasizing the value of our technique.
@@ 4.2.3 Comparison with related work
Furthermore, we investigated the potential of our approach compared to the state-of-the-art model while taking into account the amount of synthetic or real training data and the number of model parameters. We combined all the three datasets namely, GenPair , WebLI and LTIP. Table 4 compares our model’s MS-COCO benchmark performance against related methods: ITIT (Li et al., 2023b), DC (Li et al., 2023c) and SimVLM (Wang et al., 2021). Note that, in this case we provide both zero-shot and finetuning results (see Section A.2 in the Appendix). ITIT employs cycle consistency for VLM improvement, while DC uses Stable Diffusion to generate images based on human-defined heuristics, and SimVLM has one of the state-of-the-art performances on MS-COCO.
@@
Synth2 and ITIT share similar parameter counts and paired data usage, making them directly comparable. As shown in Table 4, Synth2 drastically outperforms ITIT on the CIDEr score. While DC achieves the highest raw performance, it requires significantly more parameters, relies on vast amount of real data, and the filtering procedure is heavily based on human heuristics.
@@
Synth2 strikes a balance between performance and efficiency. By leveraging text-only data, it achieves competitive results using 40 times less paired data than DC. Interestingly, despite performing weaker than SimVLM on finetuning, Synth2 has the highest in zero-shot performance among all models. This comparison highlights the trade-offs between performance, efficiency, and data requirements. Synth2 emerges as a compelling solution when resources are limited, achieving impressive results with reduced data and computational overhead.
4.2.4 Extra analysis
@@
To further analyze the boost in performance afforded by augmentation with fully synthetic data, we characterized the performance of the baseline model (i.e. trained solely on CCv2), and Synth2 which was additionally trained together with GenPair. Figure 5 visualizes the performance trends for both models. Notably, the light blue curve representing the baseline exhibits a gradual increase in performance with increasing training steps, eventually plateauing at 22.1 CIDER score. In contrast, the purple curve reveals Synth2’s steeper performance improvement, achieving comparable performance to the paired only training regime, with roughly 1/3 of the training steps. This highlights the significant data efficiency gain achieved through fully synthetic data augmentation in Synth2. Shaded regions surrounding each curve indicate standard deviation across three random seeds, demonstrating the robustness of the observed performance trends. These findings demonstrate Synth2’s ability to effectively leverage fully synthetic data, resulting in remarkable performance gains.
@@
Our study also investigated the computational efficiency of Synth2. We compared the case where captions are used to generate image embedding versus when captions are rendered into actual images. As shown in Table 5, Synth2 trained using image embedding consistently demonstrates faster performance, running at 2.08 training steps per second, whereas Synth2 trained from pixel runs at a slower pace of 1.66 steps/second. Critically, the performance on image captioning and visual question answering are not affected. This efficiency advantage stems from Synth2’s utilization of text-only data and embedding-based image generation using parallel decoding, which reduces the computational overhead associated with processing pixel-based images as done in the Synth2 model trained from pixels without affecting the performance. For further studies, e.g. comparing the performance of different image encoder backbones refer to Appendix, Table 9.
Model | All Params | Real Data | Synth Data | CIDEr (FT) | CIDEr (zs) |
ITIT | 11.2B | 3M | 110M | 103.5 | 32.1 |
DC (BLIP) | 1.7B | 5.5B | 400M | 133.1 | - |
SimVLM | 1.4B | 1.1B | 365M | 143.3 | 32.2 |
Synth2 | 632M | 10.1M | 711M | 131.3 | 35.4 |
Model | Step/Sec | MSCOCO CIDEr(zs) | VAQV2 Accuracy (ft) |
Synth2- Embeddings | 2.08 | 25.9 | 50.1 |
Synth2- Pixels | 1.66 | 26.0 | 49.9 |
@@ 5 Limitations
The two main limitations of the current work stem from the restricted quantity of fully synthetic data used and the limited exploration of text data sources. Firstly, whilst we were able to show a substantial gain when training was augmented with fully synthetic data, our experiments were limited to a relatively low quantity of data (i.e. 1M image-caption pairs). In the future, it will be important to examine whether using orders of magnitude more fully synthetic data (e.g. 700M) will result in performance gains that surpass that of using solely paired data. Also, the current work primarily explores a restricted set of text-only data sources. Further research is necessary to investigate the potential of diverse text sources, including investigating context-specific generation (e.g., medical data), which could be an exciting avenue for future exploration. By acknowledging and addressing these limitations, future research can leverage the strengths of the Synth2 approach and enhance its generalizability, robustness, and efficiency for a broader range of visual-language modeling tasks.
@@ 6 Conclusions
Our study presents a novel approach for generating synthetic image-text pairs to enhance the training of visual language models. By harnessing the capabilities of large language models and text-to-image generation, we effectively address the limitations of manual image labeling such as scalability, customization and costs. In particular, we show that the visual language model trained on our synthetic and human annotated datasets exhibits a marked improvement in image captioning tasks compared to a baseline trained exclusively on human-annotated data. Overall, this research introduces a novel technique that has the potential to transform visual language model training. Our findings highlight the value of synthetic data generation, paving the way for advancements in numerous fields where visual language understanding is crucial.
@@ Acknowledgements
We would like to thank Sander Dieleman, Ali Razavi, and Benigno Uria for their insights on VQGAN. We would also like to thank Aida Nematzadeh, Pinelopi Papalampidi, Han Zhang, Mateusz Malinowski, Valentin De Bortoli, Sahra Ghalebikesabi, Emanuele Bugliarello, Chris Knutsen, and Murray Shanahan for their in depth comments and support throughout the project.
@@ A.1.1 VQ-GAN details
Optimizer | Adam (Kingma and Ba, 2014) |
@@
Decay (Loshchilov and Hutter, 2017) Warmup steps | 10000 |
@@
A.1.2 Optimization details for Synth2
@@
Optimizer | AdamW (Loshchilov and Hutter, 2017) |
@@ A.3 Clustering details
Real | Synth | #Real Data | #Synth Data | Concentration Top-3 () | Concentration Top-5 () | Entropy () | CIDEr-COCO () | Flickr-30() | VQAV2 Acc. () | OKVQA Acc. () |
CCv2 | - | 10.1M | - | - | - | - | 22.1 | 12.7 | 29.1 | 32.4 |
CCv2 | GenPair | 10.1M | 1M | 40.5% | 57.7% | 3.81 | 25.9 | 17.3 | 31.1 | 34.0 |
CCv2+WebLI | - | 10.1M+1M | - | 58.2% | 69.8% | 3.43 | 24.4 | 14.9 | 30.6 | 33.9 |
CCv2+LTIP | - | 10.1M+1M | - | 76.5% | 83.0% | 2.92 | 23.4 | 13.8 | 30.3 | 32.9 |
A.4 Backbone Comparison
We conducted an ablation study to compare the performance of different visual backbones. For this comparison, we kept the total amount of FLOPs (both pre-training and training) and (self)supervised training data constant across models. We compared our VQ-GAN-based backbone against a contrastively pre-trained NFNet-F6 model (Brock et al., 2021) using the same loss function as CLIP (Radford et al., 2021). As shown in Table 9, the primary factor influencing performance appears to be the number of parameters, as our model with a deeper VQ-Adapter (VIT-L+) achieved results comparable to the contrastively trained model. In our study, we opted for the Synth2VIT-B backbone due to its compatibility with our hardware resources. This choice allowed us to avoid model sharding, thereby enabling faster experimental iterations.
Model | real data | #params | MSCOCO CIDEr | VAQV2 Accuracy |
Synth2 VIT-B adapt. | 690.1 | 632M | 35.3 | 37.9 |
Synth2 VIT-L+ adapt. | 690.1 | 890M | 40.9 | 42.8 |
NFNet-F6 (CLIP) | 690.1 | 920M | 41.3 | 43.4 |
A.5 Image Generator Qualitatives